OpenStudy (anonymous):

12(x + 4) = 4(1 + 3x)

rebeccaxhawaii (rebeccaxhawaii):

12x+48=4+12x 12x-12x x=0

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (imsleeziboii):

jane dont do any of what rebecca said

OpenStudy (anonymous):

why not??? @imsleeziboii

rebeccaxhawaii (rebeccaxhawaii):

lez be quiet it is correct

OpenStudy (imsleeziboii):

shes a bad person. trust me

rebeccaxhawaii (rebeccaxhawaii):

because he is all sour

OpenStudy (anonymous):

@jane.956 If you just want answers, why not use Wolfram alpha and save everyone time?

OpenStudy (imsleeziboii):

her answer is correct but she uses you

rebeccaxhawaii (rebeccaxhawaii):

how am i using her ? lol i wanna hear this

OpenStudy (imsleeziboii):

because your a feather rebecca

OpenStudy (imsleeziboii):

wtf

OpenStudy (imsleeziboii):

feather*

rebeccaxhawaii (rebeccaxhawaii):

i am thanks

OpenStudy (imsleeziboii):

im trying to say b i t c h

OpenStudy (anonymous):

ummm, I think she was just helping me..

rebeccaxhawaii (rebeccaxhawaii):

cool anyways you were rude so i didnt help you lez

OpenStudy (imsleeziboii):

i dont need your help lol

OpenStudy (imsleeziboii):

i didnt understand an you called me basically incompetent.

rebeccaxhawaii (rebeccaxhawaii):

lol k bye then but jane do you understand ?

OpenStudy (imsleeziboii):

jane dont do what she says shes retarded

candycove (candycove):

>_<

OpenStudy (anonymous):

she gave me a pretty good answer. not something a "retarded" person would do. @rebeccaxhawaii @imsleeziboii

rebeccaxhawaii (rebeccaxhawaii):

he just needs to grow up and thanks dakota

OpenStudy (imsleeziboii):

lol rebecca talk about growing up. your like 12 and you say aha whenever i post a questions

rebeccaxhawaii (rebeccaxhawaii):

im not 12 lol and you're ahella rude and get so sour so i dont help you anymore so don't act like you are too good for my help because got so many medals from helping you

OpenStudy (imsleeziboii):

i gave you 3 LMAO

OpenStudy (iluvsoccer):

Excuse me, let's not fight here. We are spamming the poor person's post. :(

OpenStudy (imsleeziboii):

Do you consider that soo many?

OpenStudy (imsleeziboii):

Sorry soccer ill be the bigger person and stop

Atsie (atsie):

@rebeccaxhawaii Rebecca honey, why don't you set the example and walk away from this post. Turn the other cheek. It tends to really slap a person in the face. :)

OpenStudy (anonymous):

considering this is none of my business and I should have never got involved, this is the last thing ima say to @imsleeziboii . clearly, you're the one that needs some growing up to do. getting mad because she's actually helping someone else? petty if you ask me. but good luck to whoever was needing and answer on this post.🤔

OpenStudy (imsleeziboii):

LOL you dont even know the story. stay out of it. i would be more than happy to attach a file of the messages she sent me